OpenStudy (anonymous):

√56 why do you have to simplify and how do you do it?

OpenStudy (anonymous):

\[\sqrt{56}\] can be simplified as \[\sqrt{2*2*2*7}\] which is \[2\sqrt{14}\]

OpenStudy (anonymous):

factorize 56, this is 7 * 2 * 2 * 2 or 7 * 2^3, So you write sqrt(7) * sqrt(2^3) = sqrt(7) * 2(sqrt(2))

OpenStudy (anonymous):

ohh I see!! sorry but why didnt you just do √7x8 = 56?